OpenStudy (veeveeniko):

What is the monthly payment on a $200,000 mortgage for 25 years at 5%?

OpenStudy (amistre64):

iknow iknow ... pick me!! :)

OpenStudy (amistre64):

how many periods are we paying for?

OpenStudy (amistre64):

12 times a year, for 25 years is how many payments to be made?

OpenStudy (veeveeniko):

300?

OpenStudy (amistre64):

and if youve already worked a solution and are trying to dbl chk, please provide your solution

OpenStudy (amistre64):

12(20+5) = 240 + 60 = 300, yes, 300 periods to pay and what is out period interest rate? 5% is a yearly rate

OpenStudy (amistre64):

*... our period rate

OpenStudy (veeveeniko):

I'm not sure

OpenStudy (amistre64):

5%/12 would be our period rate

OpenStudy (amistre64):

so our compounding rate is 1 + .05/12 per period

OpenStudy (veeveeniko):

I got 1.004

OpenStudy (amistre64):

so working the formula, mine at least P = Bk^n(1-k)/(1-k^n) = 200000(1+.05/12)^(300)(1-(1+.05/12))/(1-(1+.05/12)^(300)) and what is the resutls?

OpenStudy (veeveeniko):

1169.18

OpenStudy (amistre64):

yep, thats what i get

OpenStudy (amistre64):

that and a lot of lagtime ....